QUESTION 5 Find the derivative. + y = In(2x + 1) - e OR T T T Arial V 3 (12pt) T
Answer #1

g=lu (2x+1)- e x² = dlu (2441) d (2041) dy du d (2971) da den? da? dre drez 2 en e:22 o 2N+ 2 22 2 ne Am 2nti